Three forces of magnitude 6N,2N,3N act on a small object in directions north,south and west respectively. Find the direction and magnitude of the resultant force.If the force to move and its mass is 0.2kg.Calculate the initial acceleration.

1
Expert's answer
2022-03-13T18:46:21-0400

Given:

"F_1=\rm 6\: N"

"F_2=\rm 2\: N"

"F_3=\rm 3\: N"

The net force

"F=\sqrt{(F_1-F_2)^2+F_3^2}\\
=\sqrt{(6-2)^2+3^2}=5\:\rm N"

The direction

"\theta=\tan^{-1}\frac{4}{3}=53^\circ\; \rm N\; of\; W"

The acceleration

"a=\frac{F}{m}=\frac{5\:\rm N}{0.2\:\rm kg}=25\:\rm m/s^2"